### v1

# V-Thinker: Interactive Thinking with Images

Abstract
Empowering Large Multimodal Models (LMMs) to deeply integrate image interaction with long-horizon reasoning capabilities remains a long-standing witness in this field. Recent advances in vision-centric reasoning explore a promising “Thinking with Images” paradigm for LMMs, profoundly shifting from image-assisted reasoning to image-interactive thinking. While this milestone enables models to focus on fine-grained image regions, progress remains constrained by narrow visual tool spaces and task-specific workflow designs. To bridge this gap, we present V-Thinker, a general-purpose multimodal reasoning assistant that enables interactive, vision-centric thinking through end-to-end reinforcement learning. V-Thinker comprises two key components: (1) a Data Evolution Flywheel that automatically synthesizes, evolves, and verifies interactive reasoning datasets across three dimensions—diversity, quality, and difficulty; and (2) a Visual Progressive Training Curriculum that first aligns perception via point-level supervision, then integrates interactive reasoning through a two-stage reinforcement learning framework. Furthermore, we introduce VTBench, an expert-verified benchmark targeting vision-centric interactive reasoning tasks. Extensive experiments demonstrate that V-Thinker consistently outperforms strong LMM-based baselines in both general and interactive reasoning scenarios, providing valuable insights for advancing image-interactive reasoning applications.
1 Introduction
"The soul never thinks without an image."
— Aristotle
Humans often simplify complex problem-solving via heuristic multi-modal interactions, especially in vision‑centric reasoning [53, 67]. In challenges like geometric proofs, people interactively add auxiliary lines or sketches [19, 54], modeling visual relations to solve geometric problems more intuitively. In recent years, visual language models (LMMs) demonstrate exceptional performance in stepwise reasoning tasks, fostering expectations for their development into interactive, image-oriented thinking [70, 73, 51, 15, 21, 37, 7, 5]. However, despite producing lengthy and coherent chain‑of‑thought (CoT), these models often detach from visual grounding, leading to hallucinations [30, 28, 20]. This indicates that current visual reasoning depends more on linguistic priors than on visual perception.
To address these challenges, OpenAI’s o3 model first actively interacts with images during reasoning via visual tools (e.g., cropping, rotation), shifting the paradigm from vision‑assisted reasoning to vision‑centric thinking [39, 48]. Building on this milestone, recent efforts attempt to reproduce o3‑like interactive thinking through end‑to‑end visual reinforcement learning [81, 60]. Moreover, Thyme [78] autonomously generates executable code to render diverse visual operations on images throughout reasoning. However, the available visual actions of them are still limited and heavily depend on precise spatial localization. To further expand the visual tool space and broaden interactive thinking patterns, foundational works such as DeepSketcher [63, 72] enable LMMs to autonomously add auxiliary lines within images, explicitly modeling logical relations between visual elements. While insightful for vision‑centric reasoning, their tool designs are often tightly coupled to specific task types. Moreover, a heavy reliance on "Image2Code" pipelines for image editing struggles to accurately depict spatial relationships between visual elements and may introduce extra noise [75, 17]. This raises a fundamental research question: How can we effectively integrate image interaction into the visual chain‑of‑thought process?
In this paper, we introduce V-Thinker, a general‑purpose multimodal reasoning assistant that fosters interactive vision‑centric thinking via end‑to‑end reinforcement training. As shown in Figure 1, V-Thinker aims to simplify complex problems via autonomous interaction with images, thereby advancing the next generation of vision-centric reasoning paradigms. Specifically, V‑Thinker introduces a comprehensive vision-centric post‑training paradigm, comprising two key components: the "data evolution flywheel" and a "progressive interactive training curriculum" spanning perception to reasoning.
Data Evolution Flywheel.
An ideal vision-centric interactive reasoning model should generalize across real-world diverse tasks. To this end, V‑Thinker targets three fundamental dimensions of interactive reasoning data synthesis: (1) Diversity. We directly generate QA pairs using knowledge concepts and visual tool systems, shifting from data expansion to genuine data creation. Then, the flywheel iteratively enlarges both the concept and visual tool sets using newly synthesized data, thereby sustaining a continuous stream of diverse datasets. (2) Quality. To ensure strict quality control, we implement a coordinated calibration mechanism where a data checker rigorously screens textual, visual, and image-action dimensions while a repairer calibrates annotations across modalities, ensuring high fidelity. (3) Difficulty. Building on the preceding stages, we further introduce a progressive expansion stage that enriches the difficulty ladder by incorporating parallel and sequential extension strategies.
Through a three-stage data synthesis flywheel, we obtain a high-quality visual interactive reasoning dataset, namely V-Interaction-400K.
Visual Progressive Training Curriculum.
Following the data evolution flywheel, we design a two-stage curriculum that progressively builds robust perception and then aligns it with interactive reasoning capabilities: (1) Perception Alignment. We model the visual space through three key dimensions and synthesize perception-specific datasets to enhance the LMM’s localization capabilities, named V-Perception-40K. Furthermore, the LMM is fine-tuned on this dataset to align the localization and referencing of visual anchors. (2) Interactive Reasoning Alignment. Built upon the perceptual foundation, we implement cold-start supervised fine-tuning followed by RL within a sandboxed executor environment, enabling the LMM to autonomously generate executable code that interacts with visual elements and maintains coherent reasoning chains grounded in visual evidence.
To thoroughly evaluate LMMs’ visual-interactive reasoning capabilities, we introduce VTBench, a benchmark targeting tasks that inherently demand visual interaction. Each instance is verified by domain experts and sourced from diverse public datasets. In summary, the key contributions are as follows:
We formalize Interactive Thinking with Images and introduce V‑Thinker, an end‑to‑end multimodal reasoner that bridges visual grounding and interactive thinking through code‑driven visual tools.
We propose a Data Evolution Flywheel that automatically synthesizes, evolves, and verifies interactive reasoning datasets across three dimensions: Diversity, Quality, and Difficulty, and further release a large-scale, high-quality visual interaction dataset, V-Interaction-400K, which can also be extended to image-to-code and other vision-language tasks.
We introduce a Visual Progressive Training Curriculum that first aligns perception via point-level supervision using a high-quality dataset V-Perception-40K, and then aligns interactive reasoning through a two-stage curriculum training framework.
We introduce VTBench, an expert-reviewed benchmark featuring standardized protocols. Extensive experiments show that V-Thinker consistently outperforms mainstream LMM-based baselines on both interactive and general reasoning scenarios, providing valuable insights for advancing image-interactive reasoning.
2 Related Work
Multimodal Reasoning.
Recent advances in large language models (LLMs) and multimodal large language models (MLLMs) have significantly enhanced their reasoning capabilities across diverse tasks [29, 22, 3, 58, 23, 13, 52, 50, 2, 42, 65, 49, 41, 11, 9, 10, 12, 8, 14, 47]. Recent efforts such as MathVista [34], MathVision [56], MathVerse [76], We-Math [40], Dynamath [86], LogicVista [66], and VisuLogic [68] have introduced comprehensive benchmarks that systematically evaluate model performance across mathematical, logical, and visual reasoning scenarios. Methodologically, prior studies have improved visual–textual alignment [46, 77, 57], incorporated step-wise reasoning [85, 35], and explored RL–based optimization to strengthen multimodal reasoning [25, 74, 36, 4, 33, 1, 55, 82, 69, 43]. In particular, RL-based frameworks such as MM-Eureka [37] and Vision-R1 [25] have introduced reinforcement learning into visual reasoning, revealing new possibilities for enhancing model reasoning depth. These works have laid a foundation for the continued advancement of visual reasoning.
Thinking with Images.
Interactive visual reasoning is a long-term research vision. Early explorations, such as LLaVA-Plus [31] and Visual Sketchpad [24], pioneered this direction by enabling models to conduct visual operations during reasoning. These studies laid an important foundation for integrating interactive perception into multimodal reasoning systems. With the growing success of reinforcement learning (RL) in visual reasoning, OpenAI’s o3 [39] introduced the concept of thinking with images. Subsequent works [26, 62, 64, 83, 72, 6, 84, 71, 27, 80] have further developed this paradigm, provides a fresh perspective for the field of agentic RL. In particular, DeepEyes [81] and Thyme [79] employ RL-based training to guide models in performing executable visual tools (e.g., cropping) as part of their reasoning chain. Meanwhile, DeepSketcher [72] explores implicit visual reasoning, where models leverage abstract visual cues instead of explicit pixel-level manipulation. Building on this progress, our work advances interactive thinking with images by enabling the model to autonomously generate, execute, and iteratively refine visual code during reasoning. This opens a new perspective on integrating visual interaction into reasoning, paving the way toward more intuitive and human-like multimodal cognition.
3 Preliminary
3.1 Problem Formulation
Interactive Thinking with Images. We envision an ideal vision-centric reasoning paradigm where reasoning unfolds through progressive interaction with the image—by perceiving, modifying, and reflecting on its visual state. Building upon this idea, V-Thinker treats reasoning as a code-driven visual interaction process. At each reasoning step, the model generates a textual thought and, when necessary, a code segment that operates on the current image . The environment executes , resulting in an updated image . Formally,
where is the task query, the reasoning trajectory, and the final answer. Through this interactive process, the model reasons by generating code to modify the image and leveraging the resulting visual feedback to guide subsequent reasoning.
3.2 Rethinking on Data Synthesis Paradigm
Traditional vision-centric reasoning datasets are built upon manually defined tasks, where models act as solvers generating chain-of-thought reasoning and answers. This data synthesis paradigm limits diversity and scalability, while interactive reasoning requiring precise spatial and logical alignment remains challenging.
With the growth of modern model generation capabilities, we re-examine this paradigm from a new perspective:
In data synthesis, can models evolve from solving problems to creating problems?
To end this, we explore this paradigm from two directions.
(1) Role Transformation: from "solver" to "creator" Traditional data synthesis methods employ models as solvers to distill their reasoning trajectories, based on the core assumption that models cannot directly generate high-quality multimodal QA pairs, particularly those involving complex geometric shapes and spatial relationships. However, this approach constrains the diversity and solvability of synthesized samples.
In this work, we revisit the role of vision-language models in data synthesis, revealing that existing strong LMMs are capable of reaching the creator level in data synthesis. As shown in Figure 3, GPT-5 can directly generate Python code to render high-quality original images along with corresponding auxiliary line diagrams and reasoning trajectories.
Figure 2 shows representative rendered image comparisons. The synthesis quality far exceeds our initial expectations. The exceptional coding capabilities enable fine-grained, outstanding image editing, such as highlighting chemical elements and annotating vertical guide lines, directly creating image annotations. Therefore, we shift from the distillation paradigm, enabling innovative generation of complex problems, diagrams, and reasoning trajectories rather than passively responding to predefined tasks.
(2) Knowledge-Driven Representation. Knowledge concepts serve as anchors for human knowledge, where an ideal comprehensive knowledge system can generate reasoning data covering diverse scenarios. We therefore move beyond using seed samples as references and instead employ structured knowledge systems, where knowledge concepts serve as condensed representations of reasoning semantics, capturing diverse real-world scenarios. Knowledge specifies what to reason about, while models determine how to unfold reasoning through interaction. As shown in Figure 2, by providing only knowledge concepts and their descriptions, knowledge-driven synthesis produces problems with precise spatial alignment, coherent reasoning, and diverse visual structures, broadening the construction and evolution of reasoning data.
4 Methodology
Overview.
This section introduces V-Thinker, a general-purpose multimodal reasoning assistant that enables interactive vision-centric reasoning through end-to-end reinforcement learning. As shown in Figure 4, V-Thinker comprises two core components:
1.
Data Evolution Flywheel (§4.1): We automatically synthesize, evolve, and verify interactive reasoning datasets across three dimensions: Diversity, Quality, and Difficulty.
2.
Visual Progressive Training Curriculum (§4.2): We introduce a two-stage training framework to achieve progressive alignment from perception understanding to vision-centric interactive reasoning patterns.
Below, we delve into the specifics of our approach.
4.1 Data Evolution Flywheel
Building on the knowledge-driven paradigm outlined in § 3.2, we design the Data Evolution Flywheel, an automated, scalable, and verifiable framework for synthesizing interactive reasoning data. The framework is divided into three processes (as shown in Figure 4):
Knowledge-driven Evolution (§4.1.1) generates reasoning data through the co-evolution of knowledge and tool sets, iteratively expanding both to yield diverse problems and reasoning trajectories.
Coordinated Calibration (§4.1.2) verifies the correctness of generated questions, rendered images, and edited visual states, ensuring consistency across text, code execution, and visual outcomes.
Progressive Expansion (§4.1.3) enhances reasoning difficulty via iterative step extension and compositional knowledge integration, gradually constructing more complex and challenging reasoning chains.
4.1.1 Diversity: Knowledge-driven Evolution
Prior data synthesis methods rely on curated seed images, inherently constraining diversity. We propose that foundational knowledge concepts and visual tool system serve as the most granular anchors for synthesizing interactive data, enabling orthogonal and diverse generation. Our pipeline avoids distillation and instead instructs a strong LMM to create datasets from scratch.
Following this guideline, given an initial knowledge system derived from We-Math 2.0 [43] and a curated task-centric tool set , knowledge and tools jointly drive iterative data synthesis through a co-evolutionary loop (Algorithm 1).
At each iteration , sampled combinations from the current knowledge system are provided to the strong generator (e.g. GPT5) to construct reasoning data:
Each generated instance contains a question , the original code that renders the problem-specific image , the corresponding tool predictions , and a reasoning trajectory , where each executable code renders the visual component . Figure 3 illustrates the image rendering process from code to image.
In parallel, tool-driven generation applies the same procedure on , producing data and predicted knowledge :
Together, these complementary processes constitute a co-evolutionary loop in which and continuously generate new counterparts. The newly predicted elements are then integrated via expansion functions and :
where and filter, merge, and normalize novel elements during incorporation via BGE-based hierarchical clustering. The co-evolution mechanism among synthetic data, knowledge concepts, and visual tool systems is illustrated in the Figure 4 (left).
Through repeated execution over rounds, the system gradually enriches both and , ultimately yielding an initial dataset that serves as the foundation for subsequent calibration and expansion stages.
4.1.2 Quality: Coordinated Calibration
After obtaining the synthesized diverse dataset , strict quality control is essential. Therefore, we introduce a regulative calibration stage to ensure multi-level consistency across generated samples. This stage consists of two modules.
Checker. As shown in Figure 4 (Mid), each instance is examined by a data checking module that verifies (1) answer correctness, (2) validity of the rendered original image, and (3) coherence of intermediate visual states produced during reasoning. Only samples satisfying all three criteria are retained as valid candidates.
Repairer. For cases where the textual answer is incorrect but the rendered image is valid and intermediate visual states remain coherent, we follow the principle that a reasoning chain is fundamentally guided by its question, reconstructing the question from original and edited visual states to realign textual and visual reasoning. The reconstructed instances are re-evaluated by , and the loop repeats until inconsistencies are resolved.
Through the iterative calibration of data checking and repair, we refine into a coherent and verified dataset , which serves as the foundation for subsequent progressive expansion.
4.1.3 Difficulty: Progressive Expansion
To establish difficulty-stratified data partitions and deepen reasoning chain complexity, our intuition is to extend the CoT’s context length to achieve difficulty escalation, thereby inversely mapping more challenging QA pairs. This naturally motivates us to introduce two complementary strategies to progressively expand the difficulty: parallel and sequential extensions, constructing more challenging QA pairs:
(1) Parallel Extension. New auxiliary constructions are introduced independently of the existing ones, providing additional key observations that complement the original reasoning to reach the final answer.
(2) Sequential Extension. New auxiliary constructions are closely linked to the existing reasoning, requiring prior results or original geometric entities to define subsequent operations. For example, if the original auxiliary line is , a new line perpendicular to can be introduced to support further deductions.
Under these two strategies, a subset of verified data is sampled and provided to the expansion model, which generates extended reasoning code segments and corresponding visual states. The generated extensions are re-validated by the verification module and iteratively refined until convergence. We limit the maximum extension depth to three steps and merge all results to form the final dataset . Figure 5 display a representative sample from the synthesized dataset .
4.2 Visual Progressive Training Curriculum
For reliable vision-centric interactive reasoning, accurate perception of visual elements is crucial. While recent multimodal models demonstrate strong reasoning ability, they often struggle with fine‑grained spatial perception, failing to precisely localize points, intersections, and other anchors.
To address this, we develop a visual progressive training curriculum that aligns perception and interaction, starting with perceptual grounding through point‑level supervision (§ 4.2.1) and advancing to interactive reasoning via progressive alignment training (§ 4.3).
4.2.1 Perception Alignment
Perception Data Synthesis.
A critical aspect of perception training is synthesizing perception-specific data. Following the paradigm in Section 3.2, we model visual space via three dimensions: element relations, element count, and knowledge concepts. These aspects jointly define visual information, with style governed by knowledge and complexity by element count.
Element Relations (): We model visual elements such as points, lines, angles, and circles, based on foundational geometry principles [18], and extend this set with textual and symbolic elements for alignment with real-world scenarios. This modeling integrates both element types and their spatial relationships, such as "point on line" and "point outside circle," capturing the core geometric interactions in visual reasoning.
Element Count (): We sample the number of elements from a normal distribution to control task complexity, where represents the mean and the variance.
Knowledge Concepts (): We sample from the knowledge system to define reasoning objectives for each task.
As shown in Figure 6, each combination of these dimensions generates corresponding coordinates, which serve as visual tags representing the spatial relationships between elements. Using these coordinates, we then structure the tasks into three levels of complexity: surface-level perception, semantic-level reasoning, and integrated reasoning:
Surface-level perception: Basic tasks, such as identifying the coordinates of a specific point, for example, point .
Semantic-level reasoning: Tasks requiring geometric understanding, such as identifying the top-left vertex of a cube.
Integrated reasoning: Tasks combining perception and computation, such as finding the center of a cube based on its dimensions.
This hierarchical structure generates diverse question-answer pairs, forming the dataset , which spans all complexity levels and enhances perceptual capabilities.
Perception Training.
The model is trained using supervised fine-tuning, with the objective of minimizing the loss function:
This process enhances the model’s ability to process visual information, particularly for tasks involving point-level localization.
4.3 Interactive Reasoning Alignment
Data Selection.
For comprehensive training, we select data from the following sources:
SFT Dataset: In the Cold-Start Fine-tuning stage, we first construct the dataset to fine-tune V-Thinker with fine-grained perception capabilities, and then align to equip it with basic interactive reasoning abilities.
RL Dataset: To progressively align V-Thinker for robust interactive reasoning, we consider the following two components of the dataset: (1) Open-Source Samples: We use the We-Math 2.0 [43], MMK12 [37] and ThinkLite [59], which provide a broad range of visual reasoning data, from basic to complex, ensuring comprehensive coverage of visual reasoning tasks. (2) Targeted Sampling from : We select data from where the base model gives incorrect answers for the original image but correct answers for the edited version. A total of 3,000 instances are sampled from , ensuring a mix of simpler and more complex data, aimed at improving the model’s reasoning ability with visual states.
Cold-Start Fine-tuning
To enable the model to generate code and interact with visual elements, we apply standard supervised fine-tuning objective on the dataset , consisting of input-output pairs , where includes both the problem statement and the corresponding visual information, and represents the reasoning chain with Python code and the edited visual elements. The objective is to minimize the loss function:
This phase initially enables the model to generate code that reads and interacts with images to perform reasoning tasks.
Reinforcement Learning for Interactive Reasoning
We apply reinforcement learning (RL) to enhance the model’s interactive reasoning capabilities. Building on the Thyme [79], we use its sandbox environment to execute the model’s generated code. The model decodes reasoning tasks into executable code, which interacts with the visual elements, and the resulting outputs are fed back into the reasoning process. For optimization, we adopt Group Relative Policy Optimization (GRPO) [45], which has been shown to be effective for diverse tasks. Given an input question and a policy model , GRPO enables the reference policy to generate a group of outputs and optimizes the policy by maximizing:
where represents the importance sampling ratio for the -th token of the -th output, and denotes the advantage at time step .
Reward Design.
We follow a reward function based on the Thyme framework [79], consisting of three components: accuracy (), formatting (), and tool usage (). The total reward is defined as:
where is an indicator function ensuring that the tool usage reward is applied only when the final answer is correct and involves at least one tool. We empirically set and to balance the contributions of formatting and tool usage.
5 VTBench
To evaluate the model’s vision-centric interactive reasoning capabilities, we introduce VTBench, a benchmark designed for tasks that inherently require interaction with visual elements. The design of VTBench follows the following principles:
Specific Interactive Tasks: VTBench prioritizes tasks that inherently require interaction with visual elements, such as drawing auxiliary lines or labeling.
Expert Evaluation: Each sample is assessed by a team of five experts, who determine whether visual interaction is necessary for solving the problem. The sample is included if at least three experts agree on its necessity.
Diversity and Expansion: To ensure broad coverage, we extend the benchmark by collecting and annotating datasets from multiple sources, including both open-source benchmarks and additional tasks from public platforms, following the same annotation procedure.
5.1 Data Collection and Annotation
As shown in Figure 7, VTBench is constructed in two stages:
Sample Selection: We collect samples from open-source benchmarks: MathVista [34], MathVision [56], MathVerse [76], Dynamath [86], We-Math [40], LogicVista [66], CMM-Math [32], CharXiv [61] and ZeroBench [44], as well as additional samples gathered from public platforms. Expert evaluation is then used to determine whether interaction with visual elements is required. Each sample is included if at least three out of five experts agree on its necessity.
Interaction Annotation: We design interaction instructions based on the problem-solving chain of thought for each sample. The expert team manually generates interaction graphs in the interactive interface, which simultaneously captures perceptual coordinates. These interaction tags are then transformed into QA pairs with GPT-4.1, followed by expert validation to ensure consistency and accuracy.
5.2 Evaluation Dimensions
VTBench evaluates vision-centric interactive reasoning capabilities across three hierarchical dimensions, modeling the problem-solving process from perception to adaptive interaction during reasoning (in Figure 7):
Perception: Tasks that assess the model’s ability to recognize and interpret visual elements, such as identifying coordinates.
Instruction-Guided Interaction: Tasks where the model receives explicit instructions (e.g., drawing lines or labeling) and must interact with the visual elements to fulfill these instructions.
Interactive Reasoning: Tasks that require the model to solve reasoning tasks involving visual interaction, such as drawing auxiliary lines or modifying diagrams.
| Method | VTBench | General Reasoning | |||||||||||||
| Perception |
| GPT-4o | 2.3 | 3.7 | 38.3 | 43.8 | 68.8 | 26.3 | |||||||||
| InternVL3-78B | 10.8 | 16.0 | 43.4 | 43.1 | 64.2 | 27.7 | |||||||||
| InternVL3-8B | 10.4 | 5.4 | 36 | 29.3 | 58.8 | 24.9 | |||||||||
| LLaVA-OV-1.5-8B | 9.2 | 9.6 | 32.0 | 25.6 | 56.7 | 23.7 | |||||||||
| InternVL3-2B | 5.6 | 0.8 | 21.0 | 23.3 | 41.7 | 24.3 | |||||||||
[…]
5.3 Data Statistics and Evaluation Metrics
VTBench comprises 1,500 question-answer pairs across three task types, with 500 samples per task. It incorporates 9 open-source benchmarks across four domains (Logical Reasoning, Geometry, Algebra, Statistics), categorized into three key evaluation metrics:
For Perception Task: Considering the inconsistency in coordinate systems across different models, the model is instructed to generate Python code that draws the point at the perceived location. The generated image, compared with the annotated image, is judged by LMMs.
For Instruction-Guided Interaction Task: For tasks that require explicit instructions (e.g., drawing lines or labeling regions), the model is instructed to generate Python code to perform the required visual interaction. The result, compared with the annotated image, is judged by LMMs.
For Interactive Reasoning Task: For reasoning tasks, the model generates answers, which are then evaluated by Large Language Models (LLMs) based on correctness.
6 Experiments
6.1 Experimental Setup
Datasets.
In this paper, we use two constructed datasets for supervised fine-tuning: V-Perception-40K for perceptual alignment and V-Interaction-400K for interactive alignment. In the reinforcement learning stage, 40K samples are sampled from We-Math 2.0 [43], MMK12 [37], ThinkLite [59], and V-Interaction-400K. All datasets are curated in compliance with copyright and licensing regulations. Experiments are conducted on VTBench and three standard visual reasoning benchmarks: MathVision [56], We-Math [40], and VisuLogic [68].
Baselines.
We conduct our experiments based on the Qwen2.5-VL-7B [3] model and compare our method with two categories of baselines:
Closed-source models: For example, GPT-4o [38].
Our evaluation is performed using VLMEvalKit [16], with modifications made to the API-calling components due to network constraints.
Implementation Details.
For training, we conduct all experiments on H20 GPUs (i.e., 64 nodes with 8 GPUs per node). Both SFT stages use a learning rate of . The final SFT checkpoint initializes the RL stage, which is trained for one epoch with eight rollouts per iteration, a learning rate of , and a warm-up ratio of 0.05. For VTBench evaluation, we employ two models as judges: Qwen3-235B-A22B and Qwen3-VL-235B-A22B. Their results differ by less than 2% from GPT-4.1, validating their reliability as open-source evaluators under a unified evaluation protocol. For data construction, We employ GPT-5 as the generator and adopt Qwen3-VL-232B-A22B as the data checking module . Furthermore, GPT-4.1 is utilized as the repairer and perform progressive expansion.
6.2 Main Results
To thoroughly analyze the effectiveness of our V-Thinker in the field of interactive reasoning, we conduct experimental analyses from both quantitative and qualitative perspectives:
Quantitative Analysis.
As shown in Table 1, we compare V-Thinker with strong baseline LMMs on VTBench. V-Thinker demonstrates consistent improvements across tasks requiring vision-centric interactions, yielding the following key insights:
(1) Perception Challenges Across LMMs. Despite impressive visual reasoning abilities, perceptual alignment remains a critical bottleneck for advanced models. On VTBench, existing LMMs struggle with fine-grained visual interaction tasks, particularly in identifying spatial relationships and localizing precise points on images. While models like GPT-4o and Qwen2.5-VL excel at visual problem-solving, they underperform on tasks demanding direct visual interaction (e.g., Qwen2.5-VL: 8.8%). This reveals a substantial gap between general visual reasoning and the specific perceptual grounding required for interactive visual reasoning.
(2) Effectiveness of V-Thinker. Under identical experimental setups, V-Thinker consistently outperforms baseline LMMs across all three interactive reasoning domains, achieving an average accuracy improvement of 14.6% and maintaining over 8% gains on individual domains. Notably, it achieves over 25% performance improvement in the Instruction-Guided Interaction domain. These results underscore V-Thinker’s superior interactive thinking capability.
| Method | SFT (Per.) | SFT (Int.) | RL | MVs | WM | VS |
| ✓ | ✓ | ✓ | 29.3 | 62.8 | 26.6 | |
| ✓ | ✓ | - | 21.4 | 55.5 | 24.8 | |
| - | ✓ | ✓ | 28.3 | 62.6 | 26.1 |
Qualitative Analysis.
To further showcase V-Thinker’s beyond-expected performance in interactive thinking, we analyze three aspects: visual interaction editing, rollout sampling behavior, and fully complete cases.
(1) Visual Interaction Editing (Fig. 8): On tasks requiring visual interaction, V-Thinker-7B accurately draws squares in diagrams, fully identifying and labeling instances. For arithmetic tasks that do not strictly require interaction, the model not only produces correct answers but also proactively annotates images to clarify intermediate steps. In real-world scenarios (e.g., meal images, multi-person photos), V-Thinker likewise generates instructive visual annotations, demonstrating stable and fine-grained visual interaction editing.
(2) Rollout Sampling Analysis (Fig. 9): We quantify sampling results within a single rollout step for the same image. The results show that V-Thinker develops multi-dimensional derivative reasoning and covers a broader solution space during RL stage, reflecting stronger exploration and more diverse reasoning paths.
(3) Complete Cases (Fig. 12): To illustrate the end-to-end interactive thinking process, we present V-Thinker’s visual reasoning trajectory. During reasoning, the model autonomously generates image-editing code and immediately renders edited outputs, externalizing intermediate states and simplifying the reasoning chain—forming a think–edit loop.
Overall, these three lines of evidence jointly demonstrate that V-Thinker exhibits superior and transferable interactive reasoning capabilities.
6.3 Results on Generalized Reasoning Domain
To further validate V-Thinker’s generalization capability on general reasoning tasks, Table 1 presents its performance across three widely-used benchmarks. V-Thinker-7B demonstrates consistent improvements across three benchmarks, with stable gains in vision-centric mathematics and logical reasoning. Notably, without specific in-domain data introducing, V-Thinker achieves substantial improvements on complex multi-step reasoning tasks such as MathVision (+6.3%). These results validate that V-Thinker’s interactive reasoning paradigm generalizes effectively across diverse reasoning domains.
6.4 Ablation Study.
To explore the roles of different modules in V-Thinker, we perform an ablation study in general reasoning scenarios, as shown in Table 2. We have the following observations: (1) The performance of V-Thinker declines when any training phrase is removed, indicating that all progressive curriculum training stages are effective. (2) Removing perception alignment results in notable performance drop, demonstrating that perception capability serves as a crucial foundation for subsequent interactive reasoning. (3) Ablating RL training brings substantial performance degradation, particularly showing over 6% performance drop in Mathvision and We-Math. This indicates that RL stage is essential for exploring interactive reasoning patterns and enabling effective generalization to general reasoning scenarios.
6.5 Analysis of the Data Evolution Flywheel
Data Evolution Flywheel Effectively Expands Knowledge Systems.
Figure 10 illustrates an evolved knowledge system derived from seed concepts through the Data Evolution Flywheel, encompassing 25 distinct domains. The resulting hierarchical structure reaches a maximum depth of 7 layers and 24,767 nodes, demonstrating the flywheel’s capability to construct diverse, large-scale, and multi-level knowledge representations. This scalable architecture establishes a systematic foundation for domain-specific and cross-domain reasoning, facilitating knowledge organization and synthesis across interactive reasoning tasks.
Scaling Analysis of Evolution Iterations.
To investigate the relationship between evolution iterations and the expansion of knowledge systems and visual tools, Figure 11 quantifies the incremental growth of knowledge concepts and visual tools derived from synthesized data at each iteration.
We observe two key findings: (1) Non-linear Growth with Evolve knowledge Expansion. As evolution iterations increase, both knowledge concepts and visual tools exhibit non-linear growth without saturation. After five iterations, the system scales to approximately 50 the initial seed size. This validates the superiority of our Data Evolution Flywheel design, which fundamentally enhances data diversity by discovering orthogonal knowledge concepts and visual tools as anchors, thereby sustaining continuous flywheel momentum. (2) Impact of Initial Seed Diversity. Richer initial knowledge concepts or tool sets yield superior evolution trajectories, underscoring the critical importance of providing diverse seed knowledge concepts as foundational anchors.
7 Conclusion
In this paper, we propose V-Thinker, a general-purpose multimodal reasoning assistant that enables interactive, vision-centric thinking through end-to-end reinforcement learning. Our main contributions include: (1) formalizing Interactive Thinking with Images and developing an end-to-end framework that bridges visual grounding with code-driven interactive reasoning; (2) proposing a Data Evolution Flywheel that automatically synthesizes, evolves, and verifies datasets across diversity, quality, and difficulty dimensions; (3) introducing a Visual Progressive Training Curriculum that aligns perception and interactive reasoning through a two-stage training framework; and (4) releasing VTBench, an expert-verified benchmark for comprehensive evaluation. Extensive experiments show that V-Thinker consistently outperforms mainstream vision-language baselines, advancing the field of interactive visual reasoning and providing practical insights for future multimodal system development.
Limitation
V-Thinker explores a generalized paradigm for interactive visual reasoning, where models perform reasoning through visual interactions. However, it still has several limitations. Due to computational constraints, our current model iteration remains limited in scale and exhibits reduced capability in knowledge-intensive or domain-specific tasks. Our primary goal in this work is to establish a unified framework for interactive reasoning. In future work, we will further optimize the model and enhance its robustness and generalization across diverse reasoning domains. Moreover, through V-Thinker, we believe that as model capabilities continue to advance, both the paradigm of data construction and the upper bound of model reasoning should be re-examined and redefined. We envision future models achieving increasingly natural and human-like forms of visual interaction and reasoning.